# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ 4 Datasets
All the above datasets provide unstructured reports. Recently, Pellegrini et al. (2023) released a structured report dataset named Rad-ReStruct based on the IU-XRAY. In clinical practice, generating structured reports typically requires doctors to answer a sequence of questions (Pellegrini et al., 2023). Therefore, Pellegrini et al. (2023) designed a structured report template with a series of single- or multi-choice questions based on topic existence (e.g., Are there any diseases in the lung?), element existence (e.g., Is there an opacity in the lung?), and attributes (e.g., What is the degree?). They then integrated the IU-XRAY report data into the template using its annotated MeSH and RadLex codes.
@@
Table 1: Public datasets for medical report generation.
@@
Name Image type Report type Images Reports Patients Used by
IU X-Ray (Demner-Fushman et al., 2016) Chest X-ray Unstructured 7,470 3,955 3,955 67 works
MIMIC-CXR (Johnson et al., 2019a, b) Chest X-ray Unstructured 377,110 227,827 65,379 59 works
Chest ImaGenome (Wu et al., 2021) Chest X-ray Unstructured 242,072 217,013 4 works
COV-CTR (Li et al., 2020) Chest CT Unstructured 728 728 4 works
DeepEyeNet (Huang et al., 2021b) FA, CFP Unstructured 15,709 15,709 3 works
FFA-IR (Li et al., 2021) FFA Unstructured 1,048,584 10,790 2 works
Chinese COVID-19 CT (Liu et al., 2021e) Chest CT Unstructured 1,104 368 96 1 work
BCD2018 (Yang et al., 2021a) Breast Ultrasound Unstructured 5,349 5,349 1 work
@@
Retina ImBank (Lin et al., 2023) FP, OCT, FFA, FAF, ICG, and red-free filtered fundus images Unstructured 18,788 18,788 1 work
Retina Chinese (Lin et al., 2023) FP, FFA, and ICG Unstructured 57,498 57,498 1 work
@@
Rad-ReStruct (Pellegrini et al., 2023) Chest X-ray Structured 3,720 3,597 3,597 1 work
@@ 5 Evaluation
A precision-based metric that measures the n-gram overlapping of the generated text and ground truth text. 84 works
@@
A F1-like metric that computes a weighted harmonic mean of precision and recall based on the longest common subsequence. 75 works
@@
A F1-like metric that computes a weighted harmonic mean of unigram precision and recall. It is an extension of BLEU-1. 57 works
@@
Calculate accuracy, precision, recall, and F1 score based on observations extracted from reference reports and generated reports by automated system. 34 works
@@
Calculate precision, recall, and F1 score based on graph comparison. The ground truth and generated reports are automatically analysed to construct a sub-graph from a defined abnormality graph. 2 work
@@
Grading Radiologists need to assign a 5-point scale grade to two types of generated reports, in accordance with clinical standards. 1 work
@@ 5.1 Natural language-based evaluation metrics
Natural language evaluation metrics are from natural language processing tasks and measure the general text quality of generated reports. In the reviewed papers, the most popular metrics are BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015), which are based on n-gram matching between reference reports and generated reports. The model is deemed superior with an increased number of matches. Among them, the BLEU is the earliest and proposed a modified precision method. When evaluating the quality of radiology report generation, we typically opt for BLEU-1, BLEU-2, BLEU-3, and BLEU-4 metrics. The n in BLEU-n means the calculation is based on n-gram. The METEOR is an extension of BLEU-1 and introduces recall into evaluation. The ROUGE-L also considers precision and recall based on the longest common subsequence between reference and generated text. The CIDEr adopts the TF-IDF. The TF-IDF vectors weigh each n-gram in a sentence, and then the cosine similarity is calculated between the TF-IDF vectors of reference and generated text. When the model consistently produces the most common sentences, it can achieve notable BLEU scores. However, CIDEr can evaluate generated outputs by encouraging the appearance of important terms and punishing high-frequency vocabulary (Li et al., 2023a). CIDEr has a popular variant CIDEr-D, which introduces penalties to generate desired sentence length and remove stemming to ensure the proper usage of word forms.
@@ 5.2 Medical correctness metrics
Natural language evaluation metrics evaluate the similarity between produced reports and the ground truth, but cannot accurately measure whether the generated reports contain the required medical facts (Babar et al., 2021b; Messina et al., 2022). So medical correctness metrics are proposed to pay attention to the prediction of important medical facts. Generally, an automatic labeller is applied to extract medical facts from generated and reference reports. Then different metrics are applied to these. The mainstream metric is clinical efficacy (Liu et al., 2019; Chen et al., 2020), which initially calculates precision, recall, and F1 score, and subsequently extends to accuracy (Babar et al., 2021a; Miura et al., 2021; Moon et al., 2022; Yan, 2022; Yang et al., 2022; Selivanov et al., 2023; Yang et al., 2023) and AUC (Li et al., 2023a). Most of the works utilized CheXpert (Irvin et al., 2019) as a labeler to extract chest diseases information from ground truth reports, while seven of them (Miura et al., 2021; Yan, 2022; Wang et al., 2023a, 2024a; Dalla Serra et al., 2023b, a; Jin et al., 2024) utilized a newer labeler CheXbert (Smit et al., 2020), which has a higher performance. In addition, Pellegrini et al. (2023) generated structured reports by predicting a series of questions. They utilized macro precision, recall, and F1 score to assess all questions, along with evaluating report-level accuracy. Other metrics calculate the Hamming distance (Zhou et al., 2021) or perform graph comparison (Zhang et al., 2020) based on the extracted results.
@@ 6 Benchmark Comparison
For model comparison, it is essential to select a benchmark for an impartial evaluation. We choose the MIMIC-CXR (Johnson et al., 2019b, a)(see Table 3) dataset as a benchmark to compare the model performance for two reasons. First, according to Table 1, IU-Xray and MIMIC-CXR are the most popular datasets, but IU X-ray lacks standard training-validation-test splits, leading to less comparable results (Messina et al., 2022), while MIMIC-CXR has official training-validation-test splits. Second, MIMIC-CXR is the largest image-report dataset, which provides a broader distribution of data, facilitating testing across diverse scenarios, and reducing biases commonly encountered in smaller datasets.
@@
We endeavour to ensure equitable comparisons, but it’s important to consider the following three problems when analysing these results. First, we select the methods leveraged the official splits of the MIMIC-CXR dataset (Johnson et al., 2019b), containing 368960 images (with 222758 reports) in the training dataset, 2991 images (with 1808 reports) in the validate dataset, and 5159 images (with 3269 reports) in the test dataset. Although some papers claimed to use official splits, the total number of datasets is different. For example, the MIMIC-CXR used by Najdenkoska et al. (2022) contains 218,101 samples. These papers were excluded. In Section 3, a multitude of methodologies are highlighted. However, due to stringent screening, their absence from the comparison table does not imply inferior performance. Second, we choose BLEU, ROUGE, METEOR, CIDEr-D, precision, recall, and F1 score metrics as evaluation metric. Similar to the previous survey (Messina et al., 2022), we found that although some metrics have variants, many papers do not specify the particular version used. In that case, we assume they are consistent. Third, the generated report sections vary among different methodologies. Most articles do not explicitly specify the generated report sections, making it challenging to conduct comparative analysis.
We separate the comparison results based on the generated report section, and the completed comparison tables are provided in Table A2 in Appendix A. Table 3 shows the best and second best performances ranked for each evaluation metric. Among the 14 papers in Table 3, three of them (Song et al., 2022; Liu et al., 2023b; Jin et al., 2024) utilized multi-modality inputs. From an architectural perspective, in the encoding stage, five works utilized a pure Transformer model (Wang et al., 2022e; Kong et al., 2022; Liu et al., 2023b; Wang et al., 2023d, 2022d), another five works employed the CNN-based model (Pino et al., 2021; Song et al., 2022; Wu et al., 2022; Jia et al., 2022; Jin et al., 2024), and four works combined CNN and Transformer (Chen et al., 2021; Wang et al., 2023c, 2024a; Wu et al., 2023). In the generation stage, most papers relied on the Transformer. In terms of technical modules, six works (Wang et al., 2022e; Jin et al., 2024; Wang et al., 2023d; Wu et al., 2023; Wang et al., 2024a, 2022d) designed auxiliary tasks and five works (Song et al., 2022; Wu et al., 2022; Wang et al., 2023d; Liu et al., 2023b; Wang et al., 2022e) used contrastive learning.
@@
Table 3: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers.

### v2

@@ 6 Datasets
The findings and impressions section of the reports in the aforementioned datasets is in free-text format, which may exhibit expression ambiguity and complicate the assessment of clinical accuracy. Structured reporting presents a potential solution by employing standardized formats for documenting radiological findings and interpretations through predefined templates or checklists, such as “heart size: normal" (Nobel et al., 2022). In clinical practice, generating structured reports typically requires doctors to answer a sequence of questions (Pellegrini et al., 2023a). Pellegrini et al. (2023a) released a structured report dataset named Rad-ReStruct based on the IU-Xray. They designed a structured report template with a series of single- or multi-choice questions based on topic existence (e.g., Are there any diseases in the lung?), element existence (e.g., Is there an opacity in the lung?), and attributes (e.g., What is the degree?). They then integrated the IU-Xray report data into the template using its annotated MeSH and RadLex codes.
The training data exhibits a long-tail distribution. For example, in the IU-Xray dataset, “normal” appears 2152 times, while “degenerated” appears only once (Yang et al., 2021b). In MIMIC-CXR and CheXpert Plus, seven out of 14 diseases have positive occurrence rates below 10%, with one as low as 0.9% (pleural other in MIMIC-CXR). As a result, the model tends to favor more common words or options from the training set. Overcoming this data imbalance issue is essential.
The training and testing data for most reviewed traditional deep learning models come from the same dataset, which overlooks the importance of testing model generalization — a critical factor for clinical applications. Given the limited size of the IU-Xray dataset, several studies trained their models using other datasets and then used the entire IU-Xray dataset as an external validation set to assess the model’s generalization performance (Miura et al., 2021; Bannur et al., 2024). Moreover, clinical data exists in dynamic environments where patient populations, clinical practices, and operational processes evolve over time (Kelly et al., 2019). Therefore, for a model to be truly applicable in clinical settings, dynamic test datasets and comprehensive calibration processes are necessary to ensure the model can continue to meet requirements as data evolves.
@@
Table 2: Public datasets for medical report generation. The report format described (either free-text or structured) refers to the description of the medical condition.
@@
Name Image type Label type Images Reports Patients Used by Featured use cases
IU-Xray (Demner-Fushman et al., 2016) Frontal and lateral chest X-rays Free text report, manual/automatic keywords annotation 7,470 3,955 3,955 71 works Structured report construction
MIMIC-CXR (Johnson et al., 2019a, b) Frontal and lateral chest X-rays Free text report, automatic 14 diseases annotation 377,110 227,827 65,379 68 works Large-scale dataset, diagnosis, knowledge base
Chest ImaGenome (Wu et al., 2021a) Frontal chest X-ray Free text report, automatic objects (with locations), attributes, and relations annotation 242,072 217,013 4 works Detection, anatomy-based sentence generation, explainability, temporal change assessment
COV-CTR (Li et al., 2020) Lung CT Free text report 728 728 4 works COVID-related scenarios
DeepEyeNet (Huang et al., 2021b) FA, CFP Free text report, manual 265 diseases, and keywords annotation 15,709 15,709 3 works Diagnosis
FFA-IR (Li et al., 2021) FFA Free text report, manual retinal lesion boundary and category annotation 1,048,584 10,790 2 works Detection, explainability
COVID-19 CT (Liu et al., 2021e) Chest CT Free text report 1,104 368 96 1 work COVID-related scenarios
BCD2018 (Yang et al., 2021a) Breast Ultrasound Free text report 5,349 5,349 1 work
@@
Retina ImBank (Lin et al., 2023) FP, OCT, FFA, FAF, ICG, and red-free filtered fundus images Free text report 18,788 18,788 1 work Diverse image modalities
Retina Chinese (Lin et al., 2023) FP, FFA, and ICG Free text report 57,498 57,498 1 work
@@
Rad-ReStruct (Pellegrini et al., 2023a) Frontal and lateral chest X-rays Structured reports 3,720 3,597 3,597 1 work Structural report generation
CheXpert Plus (Chambon et al., 2024) Frontal and lateral chest X-rays Free text report, demographics, automatic 14 diseases annotation, RadGraph annotations 223,462 187,711 64,725 0 work Large-scale dataset, diagnosis, knowledge base
@@ 7 Evaluation
A precision-based metric that measures the n-gram overlapping of the generated text and ground truth text. 90 works
@@
A F1-like metric that computes a weighted harmonic mean of precision and recall based on the longest common subsequence. 81 works
@@
A F1-like metric that computes a weighted harmonic mean of unigram precision and recall. It is an extension of BLEU-1. 63 works
@@
Calculate accuracy, precision, recall, and F1 score based on observations extracted from reference reports and generated reports by automated system. 40 works
RadGraph-based metrics
(Jain et al., 2021; Delbrouck et al., 2022)
Calculate precision, recall, and F1 score based on knowledge graphs extracted from reference reports and generated reports by RadGraph series models. 5 works
@@
Calculate precision, recall, and F1 score based on graph comparison. The ground truth and generated reports are automatically analysed to construct a sub-graph from a defined abnormality graph. 2 works
@@
Grading Radiologists need to assign a 5-point scale grade to two types of generated reports, in accordance with clinical standards. 2 works
Large language model-based evaluation
GREEN (Ostmeier et al., 2024) A large model takes a reference report and a generated report as input, identifies six types of clinically significant errors, and calculates an overall score. 0 work
@@ 7.1 Natural language-based evaluation metrics
Natural language evaluation metrics are from natural language processing tasks and measure the general text quality of generated reports. In the reviewed papers, the most popular metrics are BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015), which are based on n-gram matching between reference reports and generated reports. The model is deemed superior with an increased number of matches. Among them, the BLEU is the earliest and proposed a modified precision method. When evaluating the quality of radiology report generation, we typically opt for BLEU-1, BLEU-2, BLEU-3, and BLEU-4 metrics. The n in BLEU-n means the calculation is based on n-gram. The METEOR is an extension of BLEU-1 and introduces recall into evaluation. The ROUGE-L also considers precision and recall based on the longest common subsequence between reference and generated text. The CIDEr adopts the TF-IDF. The TF-IDF vectors weigh each n-gram in a sentence, and then the cosine similarity is calculated between the TF-IDF vectors of reference and generated text. When the model consistently produces the most common sentences, it can achieve notable BLEU scores. However, CIDEr can evaluate generated outputs by encouraging the appearance of important terms and punishing high-frequency vocabulary (Li et al., 2023b). CIDEr has a popular variant CIDEr-D, which introduces penalties to generate desired sentence length and remove stemming to ensure the proper usage of word forms.
@@ 7.2 Medical correctness metrics
Natural language evaluation metrics evaluate the similarity between produced reports and the ground truth, but cannot accurately measure whether the generated reports contain the required medical facts (Babar et al., 2021b; Messina et al., 2022). Hence, medical correctness metrics are proposed to pay attention to the prediction of important medical facts. Generally, an automatic labeler is applied to extract medical facts from generated and reference reports. Then different metrics are applied to these. The mainstream metric is clinical efficacy (Liu et al., 2019; Chen et al., 2020), which initially calculates precision, recall, and F1 score, and subsequently extends to accuracy (Babar et al., 2021a; Miura et al., 2021; Moon et al., 2022; Yan, 2022; Yang et al., 2022; Selivanov et al., 2023; Yang et al., 2023) and AUC (Li et al., 2023b). Most of the works utilized CheXpert (Irvin et al., 2019) as a labeler to extract chest diseases information from ground truth reports, while seven of them (Miura et al., 2021; Yan, 2022; Wang et al., 2023a, 2024a; Dalla Serra et al., 2023b, a; Jin et al., 2024; Pellegrini et al., 2023b; Hyland et al., 2023) utilized a newer labeler CheXbert (Smit et al., 2020a), which has a higher performance. Another popular annotation tool named RadGraph extracts knowledge graphs (i.e., entities and relations) from the reports. It has an upgraded version called RadGraph-XL as mentioned in Section 3.1. Most works employed RadGraphF1 and calculated the overlap of entities and relations independently before averaging the results (Ramesh et al., 2022; Jeong et al., 2024; Hyland et al., 2023; Bannur et al., 2024). In addition, Pellegrini et al. (2023a) generated structured reports by predicting a series of questions. They utilized macro precision, recall, and F1 score to assess all questions, along with evaluating report-level accuracy. Other metrics calculate the Hamming distance (Zhou et al., 2021) or perform graph comparison (Zhang et al., 2020) based on the extracted results.
@@ 7.3 Human-based evaluation
7.4 Large language model-based evaluation
Recently, the human-like language comprehension abilities of large language models have achieved impressive performance, leading to their application in evaluation. For instance, LLaVA-Med utilizes GPT-4 to assess the helpfulness, relevance, accuracy, and level of detail of generated responses while assigning an overall score. Similarly, CheXagent provides GPT-4 with a reference report and two generated reports, enabling it to identify the superior report. However, Ostmeier et al. (2024) found that GPT-4 exhibited a low correlation with expert preferences. To address this issue, they developed a specialized model, Generative Radiology Report Evaluation and Error Notation (GREEN), to evaluate report generation. This model takes a reference report and a generated report as input, allowing it to identify and categorize errors into six categories while calculating a score. The strong correlation between this score and expert evaluations underscores its effectiveness.
@@ 8 Benchmark Comparison
For model comparison, it is essential to select a benchmark for an impartial evaluation. We choose the MIMIC-CXR (Johnson et al., 2019b, a)(see Table 4) dataset as a benchmark to compare the model performance for two reasons. First, according to Table 2, IU-Xray and MIMIC-CXR are the most popular datasets, but IU-Xray lacks standard training-validation-test splits, leading to less comparable results (Messina et al., 2022), while MIMIC-CXR has official training-validation-test splits. Second, MIMIC-CXR is the largest image-report dataset, which provides a broader distribution of data, facilitating testing across diverse scenarios, and reducing biases commonly encountered in smaller datasets. This section first introduces the process of paper selection and discusses issues related to benchmark usage. We then review the techniques employed by the top two methods at each step. Finally, we summarize the performance of large model-based approaches and compare them with traditional deep learning methods.
@@
We endeavour to ensure equitable comparisons, but it’s important to consider the following three problems when analysing these results. First, we select the methods leveraged the original official splits of the MIMIC-CXR dataset (Johnson et al., 2019b), as described in Section 6. Several methodologies highlighted in Section 3 are excluded due to their differing data utilization strategies. Given the importance of large model methods, we include their results in the comparison despite minor differences from the official test set (Wang et al., 2023e). Second, we choose BLEU, ROUGE, METEOR, CIDEr-D, precision, recall, and F1 score metrics as the evaluation metrics. Similar to the previous survey (Messina et al., 2022), we found that although some metrics have variants, many papers do not specify the particular version used. In that case, we assume they are consistent. Third, the generated report sections vary among different methodologies. Most articles do not explicitly specify the generated report sections, making it challenging to conduct comparative analysis.
We separate the comparison results based on the generated report section. The completed comparison tables are provided in Table A2 in Appendix A. Table 4 presents the top two performances for each evaluation metric. For input acquisition, three of them (Song et al., 2022; Liu et al., 2023c; Jin et al., 2024) utilized multi-modality inputs. While the ablation studies in these papers demonstrate that leveraging multi-modal data yields significant performance enhancements, particularly when real reports are retrieved and used as input (as detailed in Section 3.2) (Song et al., 2022; Jin et al., 2024), the improvements are less pronounced compared to the methods that do not employ multi-modal data. Only (Liu et al., 2023c) achieved an increase of 0.003 on METEOR, and (Jin et al., 2024) improved precision by 0.02. Thus, the utilization of multi-modal data warrants further enhancement.
@@
Table 4: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers. $\uparrow$ indicates that a higher value for this metric is better. * means large model-based methods.
@@
At the feature learning stage, for the architecture of image feature learning, four works utilized a pure Transformer model (Wang et al., 2022e; Kong et al., 2022; Liu et al., 2023c; Wang et al., 2022d), another five works employed the CNN-based model (Pino et al., 2021; Song et al., 2022; Wu et al., 2022; Jia et al., 2022; Jin et al., 2024), and four works combined CNN and Transformer (Chen et al., 2021; Wang et al., 2023c, 2024a; Wu et al., 2023b). None of the reviewed studies compared the impact of Transformer and CNN architectures on report generation tasks. In terms of training strategy and enhancement module, eight works (Wang et al., 2022e; Jin et al., 2024; Wu et al., 2023b; Wang et al., 2024a, 2022d; Wu et al., 2022; Liu et al., 2023c) leveraged auxiliary tasks and Song et al. (2022) designed contrastive attention. The ablation experiments demonstrated their effectiveness. It is apparent that using auxiliary loss for additional supervision is a widely adopted approach for enhancing report generation efficacy during the visual encoder phase. For non-image feature learning, there is a consensus on using transformer-based architectures, and for feature fusion, feature-level operations or attention are employed before inputting them into subsequent networks, lacking specific design considerations.
At the generation stage, the majority of studies employ a decoder, with only Pino et al. (2021) utilizing template-based approaches and Kong et al. (2022) adopting retrieval-based techniques. The method proposed by Kong et al. (2022) demonstrates superior performance, especially in lesion detection. It initially retrieves a substantial set of sentence candidates for subsequent selection. However, the lack of ablation studies prevents us from assessing the effectiveness of this design. Among the decoder-based methods, only two employed LSTM architecture (Wu et al., 2022; Jia et al., 2022), while the remaining methods all utilize Transformer architecture. In comparison to transformer-based approaches, LSTM-based methods exhibited inferior performance, selected primarily due to the limited number of models explicitly designed for generating the “impression” and “finding” sections. Two works utilized the modified Transformer R2Gen (Wang et al., 2023c, 2024a), and there is no significant advantage. Three studies enhanced cross-entropy loss by re-weighting (Wang et al., 2022d) or adding auxiliary loss (Wang et al., 2024a; Wu et al., 2023b), but did not yield significantly improved results. Notably, the outstanding performance of Wu et al. (2023b)’s work on the F1 metric stems from utilizing F1-based rewards in reinforcement learning, but their results on other metrics were subpar.
For large model-based methods, R2GenGPT (Wang et al., 2023e) achieved competitive results by fine-tuning only the visual encoder and adapter on the MIMIC-CXR dataset, underscoring the potential of large models in this domain. At the same time, the usage of a rich training dataset and carefully designed input-output settings enabled MAIRA-2 13B (Bannur et al., 2024) to achieve the best results in finding section generation. In addition, LLM-CXR (Lee et al., 2023) presents contrasting results on natural language metrics, with a low BLEU score (BLEU-1: 0.092) and a high CIDEr score (0.525). Disregarding differences in metric versions (e.g., CIDEr and CIDEr-D), this may suggest the model excels in keyword generation, making CIDEr a more reliable metric than BLEU due to its emphasis on keyword importance. Overall, compared with traditional deep learning-based methods, large model-based methods have significantly improved report generation performance on natural language-based metrics, achieving an improvement of more than 10% in nearly all metrics. This highlights the superior language generation capabilities of large models. However, claiming a clear advantage of large model-based methods over traditional models is difficult. First, large model-based methods typically require extensive training data. For instance, MAIRA-2 7B/13B was trained on a combination of four image-report datasets, resulting in unfair comparisons. It is essential to train traditional models on large-scale data to assess performance improvements from large model architecture or data volume. Second, advancements in medical correctness metrics are less significant, likely due to the current lack of specialized models for lesion detection in large model-based methodologies. Finally, compared to traditional models, a key advantage of large models lies in their ability to engage in interactive dialogues, which is crucial for applications such as telehealth, virtual medical consultations, and medical education (Johri et al., 2025). However, current methods fail to incorporate interactive data in large model training and lack clear evaluation protocols to assess this capability, preventing the advantage of interactive dialogue from being fully realized.
The ability of some general large models to generate reports has not been evaluated on standard benchmarks, and thus these models are not included in the table. For instance, BiomedGPT evaluated its report generation capability by randomly selecting 30 samples from the MIMIC-CXR test set using human-based evaluation. Overall, the evaluation process for large models in report generation remains inconsistent. Variations in evaluation metrics prevent a clear performance hierarchy between general and domain-specific large models. For example, miniGPT-Med outperforms the domain-specific CheXagent on BERT-Sim metric (Zhang et al., 2019) and CheXbert-Sim metric (Smit et al., 2020b), while the domain-specific MAIRA-2 surpasses the general Med-PaLM M on most metrics (e.g., ROUGE-L). However, it is difficult to compare the performance of miniGPT-Med and MAIRA-2. Recently, Zhang et al. (2024b) proposed ReXrank [^7] https://rexrank.ai/, a public leaderboard designed to evaluate AI-based report generation for chest X-ray images. This platform is essential for standardizing chest X-ray report generation benchmarks and serves as a framework for constructing benchmarks for other types of radiological images. Furthermore, an interface has been made available for users to upload their models.
In conclusion, due to the inconsistent usage of benchmarks, this section can only compare a subset of the reviewed papers. In the five stages, the integration of multimodal inputs has not demonstrated clear advantages and lacks unique designs. Current effective designs focus mainly on image feature extraction, especially through auxiliary loss. During the report generation phase, reinforcement learning effectively enhances specific metrics. However, there are currently no effective designs for loss function. Additionally, large models have shown considerable language-related capabilities, but improvements are needed to generate results with higher medical correctness. Further experiments are required to validate their superiority over traditional methods in report generation when using comparable training datasets.